OpenStudy (anonymous):

1. Factor x2 – 14x + 49 3. Factor completely. 4u2 – 20u + 25 2. What is the greatest common factor for 12xy + 16x + 20y ?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

you should use this identities (a+b)^2=a^2+2ab+b^2 (a-b)^2=a^2-2ab+b^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

for example:(x+5)^2=x^2+10x+25

OpenStudy (anonymous):

I only have 5 minutes left on my test

OpenStudy (anonymous):

for example:x^2-16x+64=(x-8)^2

OpenStudy (anonymous):

:?

OpenStudy (anonymous):

x^2-14x+49=x^2-2*7x+7^2